ABC Logistics sales are $10 million. The company spends $5 million for direct materials purchases, $2...

  1. ABC Logistics sales are $10 million. The company spends $5 million for direct materials purchases, $2 million for direct labor purchases. Overhead is $2 million and profit is $1 million. Direct labor and direct materials vary with sales, but overhead (fixed costs) does not. The company wants to double its profit.
    1. By how much should the firm increase annual sales?
    2. By how much should the firm decrease material costs?
    3. By how much should the firm decrease labor costs?
Answer #1

a.

At present the direct material is 50% of the annual sales and the direct labor is 20% of the annual sales. If the an annual sales of X gives a profit of 2 million then we can equate it as

X – 0.5X – 0.2X – 2 = 2

Solving this, we get

X = 4/0.3 = 13.33 million.

If the firm increases the annual sales by 3.33 million then they can double the profit.

b.

Right now the material cost is 5 million and we need to increase our profit by 1 million. Thus if we reduce the material cost by 1 million to 4 million then our profit will be 10 – 4 – 2 – 2 = 2 million.

c.

Right now the labor cost is 2 million and the target is to make additional 1 million. If we want to reduce labor cost to double the profit, we should reduce labor cost to 1 million. The profit will be 10 – 5 – 1 – 2 = 2 million.